Alice has $225,000 in savings in an investment account. She is able to earn an APR of 4.5% in her investment account (which has monthly compounding). Ignore tax.

She decided to buy a house of take out a 30-year mortgage. The mortgage will have an APR of 4.5% with monthly compounding and monthly payments. The bank requires at least a 20% house value down payment. She decides to buy the largest possible house given her savings.

What is her monthly mortgage payment?
Answer #1

Amount of loan = 225000*80/20 (Since the savings will be used to pay 20% down payment and balance 80% of value will be the loan amount)

= $900,000

Using financial calculator
Input: N = 30*12=360

PV = -900000

I/Y = 4.5/12

Solve for PMT as 4560.17

The monthly mortgage payment is

$4,560.17